OpenStudy (anonymous):

solve the system of equations by elimination 3x+6y=5 -3x-3y=7

jimthompson5910 (jim_thompson5910):

notice you have +3x in equation 1 and -3x in equation 2

jimthompson5910 (jim_thompson5910):

they add to 3x + (-3x) = 3x - 3x = 0x = 0 so effectively adding them gives you 0 (which goes away)

jimthompson5910 (jim_thompson5910):

this means that if you added both equations, you would get 3x+6y=5 -3x-3y=7 ----------- 0x + 3y = 12 so we know that 3y = 12, which means y = ??? once you figure out y, you can solve for x

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

waiiiit I needed help solving x

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

but somehow I got 3x=-19 for x and it doesn't go

jimthompson5910 (jim_thompson5910):

good y = 4 is correct

jimthompson5910 (jim_thompson5910):

3x+6y=5 3x+6*4=5 ... plug in y = 4 3x + 24 = 5 3x = 5 - 24 3x = -19 x = -19/3 you can leave it as a fraction or convert to decimal form (it will be approximate though)

jimthompson5910 (jim_thompson5910):

OR you can pick on the second equation -3x-3y=7 -3x-3*4=7 -3x - 12 = 7 -3x = 7 + 12 -3x = 19 x = 19/(-3) x = -19/3 either way, you'll get x = -19/3

OpenStudy (anonymous):

okay I thought I did something wrong

jimthompson5910 (jim_thompson5910):

nope you did it perfectly

jimthompson5910 (jim_thompson5910):

it's perfectly possible to get a fraction as an answer

OpenStudy (anonymous):

ahh okay then thanks again

jimthompson5910 (jim_thompson5910):

sure thing